OpenStudy (anonymous):

**i will give a metal to anyone can answer this correct** Evaluate the function for the given value of x. f(x) = -6x2 - 4x - 4 , x = - 2

OpenStudy (anonymous):

-6x2 = 24 -4x = 8 -4 = 28 assuming you mean -6 * x * 2

OpenStudy (anonymous):

f(x)=-6x^2-4x-4 so, f(-2)=-6(-2)^2-4(-2)-4 =-6(4)+8-4 =-24+4 =-20